OpenStudy (anonymous):

A man invests $5,200, part at 4% and the balance at 3%. If his total income for the two investments is $194, how much money did he invest at each rate?

OpenStudy (anonymous):

Interest = Investment * Rate X = amount invested at 4% $5,200 - X = amount invested at 3% 194 = 0.04X + 0.03(5,200 - X) 194 = 0.04X + 156 - 0.03X 38 = 0.01X X = 3800 Answer: $3,800 is invested at 4% and $1200 is invested at 3% Check: $3,800 * 4% = $152 $1,200 * 3% = $36 $152 + $36 = $192